OpenStudy (anonymous):

4x(x − 2) −3x(x + 1) = −18

OpenStudy (mayankdevnani):

4x^2-8x-3x^2-3x=-18

OpenStudy (mayankdevnani):

can you solve it

OpenStudy (mayankdevnani):

@warsop87 can you

OpenStudy (anonymous):

Give me a sec

OpenStudy (mayankdevnani):

x^2-11x=-18

OpenStudy (mayankdevnani):

there can you

OpenStudy (mayankdevnani):

can you solve it? @warsop87

OpenStudy (anonymous):

Im trying I suck at math and not as good as you that I can see

OpenStudy (mayankdevnani):

ok so m solve your question can i

OpenStudy (anonymous):

Yes please

OpenStudy (mayankdevnani):

x^2-11x=-18

OpenStudy (mayankdevnani):

then x^2-x=-18/11

OpenStudy (anonymous):

Thank you! ;)

OpenStudy (mayankdevnani):

ok